Please help me Find the value of 'a' so the equation 4x-3a=2(x+a) has the solution x=10
Answer by Alan3354(69443)

Find the value of 'a' so the equation 4x-3a=2(x+a) has the solution x=10
---------
Sub 10 for x, solve for a.
4x-3a=2(x+a)
4*10 - 3a = 2(10-a)
40 - 3a = 20 - 2a
a = 20